OpenStudy (anonymous):

How do you express Sin315 as a function

OpenStudy (ikram002p):

sin 315 is not a function :O its a value / constant value how ever sin x is a function of variable x

OpenStudy (ikram002p):

and since f(x) = sin x we can say that f(315) =sin 315 how ever i think u need what slash eplained , not what u asked ffor

OpenStudy (anonymous):

sin(315)=sin(360-45); sin(360-x)=-sin(x); therefore, sin 315 = -sin (45)= -0.707